17.5 L N2 at 25 °C and 125 kPa and 39.5 L O2 at 25 °C and 125 kPa were transferred to a tank with a volume of 7.00 L. What is the total pressure at 35 °C?

Answer in kPa
Answer #1

Answer:

Step 1: Explanation

We know the ideal gas equation PV=nRT

where, R=universal gas constant, T=temperature , P=pressure, V=volume, n=moles.

Step 2: calculate the moles for N2 and O2 each

Moles of N2 = pv/RT = (125 kPa) x (17.5 L) / ((8.3144621 L kPa/K mol) x (25+ 273.15 K)) =0.882 mol of N2

Moles of O2 = (125 kPa) x (39.5 L) / ((8.3144621 L kPa/K mol) x (25+273.15 K)) =1.992 mol of O2

Step 3: Calculation of  the total pressure

Total moles = ( 0.882+1.992 ) moles = 2.874 mol

volume = 7 L and Temperature = (35 +273.15)K = 308.15 K

By using the ideal gas equation

P = nRT / V

P = (2.874 mol) x (8.3144621 L kPa/K mol) x (308.15 K) / ( 7 L) = 1052  kPa

Hence , the total pressure at 35 °C is 1052 kPa